Calculate the pH of a solution with a [OH -] concentration of 2.3 X 10 -7 M.
1
Expert's answer
2015-01-15T09:33:06-0500
First we calculate the pOH. pOH = -log(2.3 X 10 -7) = 6.64
pH + pOH = 14, so pH = 14-6.64 = 7.36